OpenStudy (anonymous):

(6+6i)(-3+3i) put in trigonometric form

OpenStudy (anonymous):

before or after you multiply?

OpenStudy (anonymous):

either way i guess.

OpenStudy (anonymous):

if you multiply you get -36

OpenStudy (anonymous):

which in trig form would be \[36(\cos(\pi)+i\sin(\pi))\]

OpenStudy (anonymous):

how did you get -36?

OpenStudy (anonymous):

my guess is you are supposed to put them in trig form first

OpenStudy (anonymous):

well no, you got the right answer, i just want to know how to do it

OpenStudy (anonymous):

\[(a+bi)(c+di)=(ac-bd)+(ad+bc)i\] so \[(6+6i)(-3+3i)=(-18-18)+(18-18)i=-36\]

OpenStudy (anonymous):

trig form of \(6+6i\) is \(6\sqrt{2}(\cos(\frac{\pi}{4})+i\sin(\frac{\pi}{4}))\)

OpenStudy (anonymous):

angle should be clear because you are over 6 and up 6, and \(r=\sqrt{6^2+6^2}=\sqrt{2\times 6^2}=6\sqrt{2}\)

OpenStudy (anonymous):

similarly \(-3+3i=3\sqrt{2}(\cos(\frac{3\pi}{4})+i\sin(\frac{3\pi}{4}))\)

OpenStudy (anonymous):

okay, i get it thank you(:

OpenStudy (anonymous):

yw